OpenStudy (anonymous):

HELP WITH ACT MATH REVIEW QUESTION IMPORTANT: the depth of a pond is 180 cm and is being reduced by 1 cm per week. the depth of a second pond is 160 cm and is being reduced by 1/2 cm per week. If the depths of both ponds continue to be reduced at these constant rates, in about how many weeks will the ponds have the same depth?

OpenStudy (anonymous):

@Vocaloid @mathstudent55 @Nnesha @BAdhi @mathmate @LexiLuvv2431 @lily1021 @RiOT @geny55 @emogirl100 @Elizabeth_46 @flatouthero @LegendaryNikki @Princessroniya

OpenStudy (anonymous):

Yo

Vocaloid (vocaloid):

write equations for both ponds depths pond 1: d = 180-x pond 2: d= 160-(0.5x) where x = number of weeks and d = depth set them equal to each other and solve for x 180-x = 160-(0.5x)

OpenStudy (anonymous):

@Vocaloid why? if i don't no one helps me with my problem

Vocaloid (vocaloid):

tagging = okay tagging a ton of people at a time = generally not ok

OpenStudy (anonymous):

but why? just curious

Vocaloid (vocaloid):

everyone who gets tagged gets a bunch of notifications

OpenStudy (anonymous):

okay sorry

Vocaloid (vocaloid):

anyway, back to the problem...

OpenStudy (anonymous):

also could you help me try to actually solve for x please?

Vocaloid (vocaloid):

180-x = 160-(0.5x) start by adding x to both sides

OpenStudy (anonymous):

so would i get 180-.5x=160

Vocaloid (vocaloid):

yes, keep going now subtract 180 from both sides

OpenStudy (anonymous):

I'm just wondering i thought i would add 180 plus 160

Vocaloid (vocaloid):

nope, they're not on the same side of the equal sign so we subtract 180 - 0.5x = 160 subtract 180 from both sides, what do you get?

OpenStudy (anonymous):

-.5x=-20

Vocaloid (vocaloid):

right, now we divide both sides by (-0.5)

OpenStudy (anonymous):

its 40! thank you so much for the help :-)